# Citation Sentiment Reflects Multiscale Sociocultural Norms

### v1

@@ II.4 Citation Sentiment Tracks Country Differences in Individualism and Power
Figure 6: Critical sentiment varies across countries. (A) Mean critical sentiment across countries toward collaborators versus non-collaborators. (B) Mean critical sentiment for countries as a function of the degree to which people accept the unequal distribution of power, measured as the power distance [61]. (C) Mean critical sentiment for countries as a function of individualism; individualism measures the degree to which people believe that the interests of the individual should take precedence over the interests of a social group[61]. (D) Relationship between the mean and bias in critical sentiment across countries. (E) Bias in critical sentiment across countries as a function of power distance. (F) Bias in critical sentiment across countries as a function of individualism.
@@
Accordingly, we examine power distance and individualism—two key dimensions of modern culture [61, 74]—to assess country-level differences in citation sentiment. Power distance refers to the degree to which people accept the unequal distribution of power. Individualism refers to the degree to which people believe that the interests of the individual should have precedence over the interests of a social group, and to which people oppose external interference upon individual interests by society or institutions such as the government [61]. We observe a significant negative correlation between critical sentiment and power distance (Fig. 6B; linear regression adjusted $R^{2}=0.737$; F-test $F(1,21)=62.50$, $p<0.0001$), such that countries whose people accept the unequal distribution of power tend to evince less critical citation sentiment. The effect holds for both non-collaborators (adjusted $R^{2}=0.756$, $F(1,21)=69.24$, $p<0.0001$) and collaborators (adjusted $R^{2}=0.427$, $F(1,21)=17.37$, $p<0.001$). Further, we observe a significant positive trend between critical sentiment and individualism (Fig. 6C; linear regression adjusted $R^{2}=0.552$; F-test $F(1,21)=28.13$, $p<0.0001$), such that countries whose people privilege individual interests over group interests tend to evince more critical citation sentiment. This effect again holds for both non-collaborators (adjusted $R^{2}=0.588$, $F(1,21)=32.41$, $p<0.0001$) and collaborators (adjusted $R^{2}=0.213$, F-test $F(1,21)=6.96$, $p<0.05$). The overall trends between favorable sentiment and power distance (Fig. 7B, grey) or individualism (Fig. 7C, grey) are similar but weaker in size compared to those of critical sentiment. Interestingly, unlike critical sentiment, an opposite trend appears between favorable sentiment towards collaborators and power-distance (Fig. 7B, cyan, adjusted $R^{2}=0.197$, F-test $F(1,21)=6.40$, $p<0.05$). The bias in critical sentiment towards non-collaborators as opposed to collaborators (Fig. 6D) does not track individualism or power-distance (Fig. 6E-F), but favorable sentiment does (Fig. 7E, adjusted $R^{2}=0.300$, F-test $F(1,21)=10.42$, $p<0.01$ for power-distance; Fig. 7F, adjusted $R^{2}=0.240$, F-test $F(1,21)=7.96$, $p<0.05$ for individualism). While there is a trend between average favorable sentiment expression and the bias towards collaborators, this trend may not be robust due to the large number of points that do not follow it (Fig. 7D). These results suggest that norms around social hierarchy may influence citation sentiment, resulting in a stronger expression of differences and criticisms in low-power and high-individualism cultures.
@@
Figure 7: Favorable sentiment varies across countries. (A) Mean favorable sentiment across countries toward collaborators versus non-collaborators. (B) Mean favorable sentiment for countries as a function of the degree to which people accept the unequal distribution of power, measured as the power distance [61]. (C) Mean favorable sentiment for countries as a function of individualism; individualism measures the degree to which people believe that the interests of the individual should take precedence over the interests of a social group[61]. (D) Relationship between the mean and bias in favorable sentiment across countries. (E) Bias in favorable sentiment across countries as a function of power distance. (F) Bias in favorable sentiment across countries as a function of individualism.
@@ II.5 Citation Sentiment and Gender
We observe that men display $2.9\%$ more critical sentiment than women (Fig. 8A; two-sided Welch’s t-test: $t=51.02$, $p<0.0001$, $df=1773$) and more favorable sentiment than woman (Fig. 8A; two-sided Welch’s t-test: $t=92.69$, $p<0.0001$, $df=1750$). This pattern of results indicates that men write with more sentiment overall than women. Interestingly, men also display $0.6\%$ less favorable sentiment bias than women (two-sided Welch’s t-test: $t=7.73$, $p<0.0001$, $df=1738$; Fig. 8A), and $2.7\%$ less critical sentiment bias than women (two-sided Welch’s t-test: $t=-15.23$, $p<0.0001$, $df=1813$). This pattern of results indicates that men more evenly distribute their sentiment across both collaborators and non-collaborators compared to women. Because gender ratios vary by discipline [85, 86, 59], we next examined whether citation sentiment tracked the proportion of men in a discipline. We observe that the proportion of men is not correlated with critical sentiment (Fig. 8B; linear regression adjusted $R^{2}=0.049$; F-test $F(1,25)=2.35$, $p=0.14$) but it is positively correlated with favorable sentiment towards collaborators (Fig. 8C; adjusted $R^{2}=0.483$; F-test $F(1,25)=25.33$, $p<0.0001$) and towards others in general (Fig. 8C; adjusted $R^{2}=0.120$; F-test $F(1,25)=4.54$, $p<0.05$). Generally, these results indicate that gender interacts with both discipline and collaboration status in explaining citation sentiment.
@@ III.3 Relations of culture across countries
Although scientists live and work within an explicitly scholarly social milieu—comprised of collaborative relationships and disciplinary identities—science and its practitioners also exist within other cultural structures outside of the academy. For example, research priorities and funding levels differ significantly across countries [120, 121]. Further, social structures of prestige and marginalization impact what science is done and who gets to do it [122, 24], often with detrimental outcomes for discovery and innovation [123, 124]. Here we asked whether the well-known cultural differences in academic writing extend to citation practices, and whether the sentiment of those citations tracks cultural norms around status, hierarchy, and power. We found that countries characterized by individualism displayed greater citation sentiment (both critical and favorable) to non-collaborators, indicating a more pervasive use of affect in scientific writing. By contrast, countries characterized by collectivism displayed less citation sentiment, utilizing a more neutral writing style. We further found that countries whose people tended to accept the unequal distribution of power (as quantified by the power distance) displayed less citation sentiment (both critical and favorable) to non-collaborators. It is possible that citation sentiment is used in science to uphold or dismantle existing structures of status, hierarchy, and dominance whereas the lack of sentiment indicates an acceptance of power and a focus on collective goals. More broadly, our results suggest that the non-scholarly cultural systems that we as scientists are embedded within play a key role in how scholars practice science individually.
III.4 Methodological Considerations
@@
Several methodological considerations are pertinent to our study. First, our sentiment analysis method captures both evaluative judgments and expressed commonalities and differences between the citing and the cited works. Sentiment can be defined in multiple ways [125], and the findings must be interpreted in the context of the definition used. Second, because we have limited amount of data for each individual author, we cannot predict how a single citer might behave. Instead, our results characterize trends for a group or groups of people. Third, in seeking to understand how scholars cited the work of others with different sorts of affect, it was important to ensure that differences in scholarly content were not driving our findings. Hence, we developed and employed a null model that accounts for differences in content between citing and cited papers, thereby ensuring a clean representation of sentiment. Fourth, studies of culture are challenged by the complexity of the concept, and prior work has underscored the need to think about multiple dimensions and scales of culture [65, 66]. Here we acknowledge that complexity by examining cultural factors that span from small to large scales, and concretize our investigation by using specific variables related to social practices and beliefs. Fifth, prior work has found that criticism and intellectual differences in citations are rare (from $0.8\%$ [126] to $15\%$ [127]). To ensure a reliable signal, we employed a large database of $627108$ citation sentences and a powerful and sensitive sentiment detection method using a large language model. Sixth, our results are limited to the field of neuroscience, and it is not yet clear whether they are pertinent to other fields. It seems likely that there could be links to other young disciplines, that have common factors like a relative youth which could allow for the emergence of still-malleable subdisciplines, a marked interdisciplinarity drawing on older fields, and a diversity of opinions, methods, and experimental approaches [36, 37]. It would be interesting to contrast this work to other disciplines of varying ages.
@@ V.1 Citation and Author Data
We use PubMed Central and the Web of Science API to build our datasets for analyses. The PubMed Central (PMC) Open Access Subset (OA) includes millions of research articles and pre-prints alongside relevant metadata, such as article type, venue of publication, and institutional information [129]. We gather all English language research and review articles from PMC OA relating to neuroscience published between $1998$ and $2023$ and cited at least one article. In parallel, using the Web of Science, we compile a list of journals classified under the “neurosciences” category with an impact factor of at least $3$ in the year $2022$. We manually add three more venues creating a list of $181$ journals in which the gathered articles are published; the full list is available in the Supplement. Each article in the PMC OA subset is stored in an extensible markup language (XML) file and contains tags delineating its several components. For instance, the <article-title> tag encloses an article’s title, <abstract> encloses the abstract, and <xref> encloses a citation in the body of the text. Using these and other relevant tags, we extract each article’s title, list of authors, affiliations (at the departmental, institutional, and country levels), and all sentences containing citations to other works. We obtain author h-indices from the Web of Science. We use Gender API to infer gender from the names of the last authors, requiring the underlying model to be at least $70$% confident in its prediction [130]. The resulting dataset contains $627108$ citation sentences, $383292$ unique authors, $56034$ unique last authors; from the data, $27$ departments and $23$ countries and regions are included in the analyses. Additionally, we build a collaboration network where authors are nodes and edges correspond to co-authorship. From this network, we define collaboration distance as the length of the shortest path between two individuals.
@@ V.2 Measuring Sentiment
We assess citation sentiment using GPT-3.5-Turbo-1106, a large language model (LLM) by OpenAI. LLMs are trained on a large corpus of text data and can generate language across various domains. Through their exposure during training to diverse contexts, they can recognize subjective elements of human language, such as tone [131], irony [132, 133], sarcasm [134], and implied meaning [135]. Here, we prompt GPT-3.5-Turbo-1106 to evaluate a citation as negative, neutral, or positive; prompt text supplied to the LLM is available in the Supplement. Each entry in the citation data contains a single sentence with possibly multiple citations. Consider, for instance, the following sentence with two citations: “Lorem ipsum odor amet [citation 1], consectetuer adipiscing elit [citation 2].” This text results in two entries for the LLM to judge: (1) “Lorem ipsum odor amet [citation 1], consectetuer adipiscing elit”, and (2) “Lorem ipsum odor amet, consectetuer adipiscing elit [citation 2].” Using the OpenAI API, we pass each entry to the model separately for sentiment analysis. In practice, each citation location in a sentence is replaced by the Unicode character with encoding U+272A (see Supplement for full model prompt). It is common for an article to cite another article multiple times in different locations in the text. We aggregate overall sentiment through the following order of precedence: negative sentiment takes precedence over positive sentiment, and both take precedence over neutral sentiment. Here, we assume negative sentiment from a citing article overrides any positive or neutral citations[136, 137].
@@
However, before using the model in this manner, we perform an additional fine-tuning step. LLMs can perform various tasks through zero-shot prompting, meaning they do not generally require more training. Supplying a well-chosen prompt is often sufficient. However, a small amount of domain-specific finetuning can improve performance and help generate well-formatted outputs for easier parsing. The finetuning process entails adapting a large language model trained for a general use case to a task-specific use case, in this case, assessing citation sentiment. With the help of $5$ volunteer researchers, we manually annotate $300$ randomly chosen citations as negative ($-1$), neutral (00), or positive ($1$). The finetuning dataset comprises these citations and their annotator-assigned labels. If the average sentiment value is less than or equal to $-0.4$, we assign a negative label; if it is greater than or equal to $0.4$, we assign a positive label. All citations with values between the two are assumed to be neutral. Finetuning GPT-3.5-Turbo-1106 with this data results in an improvement of $28\%$ over the baseline (see Supplement).
@@
We use a null model to benchmark observed sentiment values against those that might emerge by random chance. For instance, we find that papers with similar content are likely to involve negative citations, thereby increasing the likelihood that such a citation in the data is an outcome of the similarity of work and not an underlying bias of interest. Similar effects are associated with article type (review vs. research) and the number of times an article references another. A review article tends to be more neutral in sentiment compared to a research article, and an article that cites another article more frequently tends to be more negative. We use these factors to measure the expected probability of observing a negative, neutral, or positive citation in the data. First, relating to shared content, we create three bins for high, medium, and low similarity, where similarity is measured as the cosine distance between the vector embeddings of article titles. The embeddings are generated using OpenAI’s text-embedding-3-small model. Additionally, $15$ bins are associated with citation frequency and $2$ with article type. Initially, for each of the $90$ ($3\times 15\times 2$) resulting combinations, we perform maximum likelihood estimation to compute the probability that a citation in the data belongs to a given group. However, if a bin has fewer than $500$ samples, we use the marginal distribution over title similarity to create a larger bin. In these instances, we choose to ignore the impact of title similarity since it has the weakest effect on citation valence compared to article type and citation frequency. Further, when a citing-cited pair has a frequency greater than $15$, we use the corresponding bin for $15$ citations instead. We use the probabilities underlying the $54$ resulting bins to draw realizations of null sentiment values for given citing-cited pairs. These realizations serve as the basis of comparison throughout our analyses.
@@
Consider a set of citations corresponding to a broader category of interest, such as citations between collaborators or citations between authors with an h-index difference of $2$. Given such a set, we can measure the ratio of positive, neutral, and negative citations within it. However, if the set is small in size, the estimated ratios may not accurately represent the true underlying distribution of sentiments for the group. To mitigate this concern, we apply bootstrapping. We create $1000$ resampled sets by randomly sampling from the original, each time with replacement, thereby allowing us to estimate a distribution of ratios for each type of sentiment. Bootstrapping in this manner helps to construct robust distributions and derive more accurate measures of variability. Simultaneously, for each of the $1000$ resampled sets for a given sentiment type, we evaluate expected ratios considering the null model. Our measure of interest is the percentage difference in the empirical ratio relative to the ratio derived from the null model. In all figures, we plot mean values for the normalized empirical ratios and estimate error bars using standard deviations.

### v2

@@ II.4 Citation Sentiment Tracks Country Differences in Individualism and Power
Figure 6: Critical sentiment varies across countries. (A) Mean critical sentiment across countries toward collaborators versus non-collaborators. (B) Mean critical sentiment for countries as a function of the degree to which people accept the unequal distribution of power, measured as the power-distance [66]. (C) Mean critical sentiment for countries as a function of individualism; individualism measures the degree to which people believe that the interests of the individual should take precedence over the interests of a social group [66]. (D) Relationship between the mean and bias in critical sentiment across countries. (E) Bias in critical sentiment across countries as a function of power-distance. (F) Bias in critical sentiment across countries as a function of individualism.
@@
Accordingly, we examine power-distance and individualism—two key dimensions of modern culture [66, 79]—to assess country-level differences in citation sentiment. Power-distance refers to the degree to which people accept the unequal distribution of power. Individualism refers to the degree to which people believe that the interests of the individual should have precedence over the interests of a social group, and to which people oppose external interference upon individual interests by society or institutions such as the government [66]. We observe a significant negative correlation between critical sentiment and power-distance (Fig. 6B; linear regression slope $s=-0.69$, $95\%$ CI $(-0.87,-0.50)$, adjusted $R^{2}=0.729$; one-sided F-test $F(1,21)=60.06$, $p<0.0001$; $n=23$; all country-level results share the same sample size), such that countries whose people accept the unequal distribution of power tend to evince less critical citation sentiment. The effect holds for both non-collaborators ($s=-0.73$, $95\%$ CI $(-0.91,-0.55)$, $R^{2}=0.760$, $F(1,21)=70.65$, $p<0.0001$) and collaborators ($s=-0.63$, $95\%$ CI $(-0.95,-0.30)$, $R^{2}=0.407$, $F(1,21)=16.09$, $p<0.001$). Further, we observe a significant positive trend between critical sentiment and individualism (Fig. 6C; linear regression slope $s=0.39$, $95\%$ CI $(0.24,0.54)$, adjusted $R^{2}=0.550$; one-sided F-test $F(1,21)=27.87$, $p<0.0001$), such that countries whose people privilege individual interests over group interests tend to evince more critical citation sentiment. This effect again holds for both non-collaborators ($s=0.42$, $95\%$ CI $(0.26,0.57)$, $R^{2}=0.587$, $F(1,21)=32.23$, $p<0.0001$) and collaborators ($s=0.30$, $95\%$ CI $(0.05,0.54)$, $R^{2}=0.198$, one-sided F-test $F(1,21)=6.44$, $p<0.05$). The overall trends between favorable sentiment and power-distance (Fig. 7B, grey) or individualism (Fig. 7C, grey) are similar but weaker in size compared to those of critical sentiment. Interestingly, unlike critical sentiment, an opposite trend appears between favorable sentiment towards collaborators and power-distance (Fig. 7B, cyan, slope $s=0.18$, $95\%$ CI $(0.03,0.34)$, adjusted $R^{2}=0.180$, one-sided F-test $F(1,21)=5.83$, $p<0.05$). The bias in critical sentiment towards non-collaborators as opposed to collaborators (Fig. 6D) does not track individualism or power-distance (Fig. 6E-F), but favorable sentiment does (Fig. 7E, slope $s=-0.26$, $95\%$ CI $(-0.42,-0.09)$, adjusted $R^{2}=0.287$, one-sided F-test $F(1,21)=9.88$, $p<0.01$ for power-distance; Fig. 7F, slope $s=0.16$, $95\%$ CI $(0.04,0.28)$, adjusted $R^{2}=0.237$, one-sided F-test $F(1,21)=7.84$, $p<0.05$ for individualism). We observe a weak trend between average favorable sentiment expression and the bias towards collaborators (Fig. 7D). These results suggest that norms around social hierarchy may influence citation sentiment, resulting in a stronger expression of differences and criticisms in low-power and high-individualism cultures.
@@
Figure 7: Favorable sentiment varies across countries. (A) Mean favorable sentiment across countries toward collaborators versus non-collaborators. (B) Mean favorable sentiment for countries as a function of the degree to which people accept the unequal distribution of power, measured as the power-distance [66]. (C) Mean favorable sentiment for countries as a function of individualism; individualism measures the degree to which people believe that the interests of the individual should take precedence over the interests of a social group [66]. (D) Relationship between the mean and bias in favorable sentiment across countries. (E) Bias in favorable sentiment across countries as a function of power-distance. (F) Bias in favorable sentiment across countries as a function of individualism.
@@ II.5 Citation Sentiment and Gender
Compared to the null model, we observe that men display $2.89\%$ more critical sentiment than women (Fig. 8A; two-sided Welch’s t-test: $t=52.55$, $p<0.0001$, $df=1851$) and $1.72\%$ more favorable sentiment than woman (Fig. 8A; two-sided Welch’s t-test: $t=95.40$, $p<0.0001$, $df=1787$). This pattern of results indicates that men write with more sentiment overall than women. Interestingly, men also display $0.63\%$ less favorable sentiment bias than women (two-sided Welch’s t-test: $t=8.46$, $p<0.0001$, $df=1741$; Fig. 8A), and $2.63\%$ less critical sentiment bias than women (two-sided Welch’s t-test: $t=-14.66$, $p<0.0001$, $df=1808$). This pattern of results indicates that men more evenly distribute their sentiment across both collaborators and non-collaborators compared to women.
Because gender ratios vary by discipline [90, 91, 64], we next examined whether citation sentiment tracked the proportion of men in a discipline. We observe that the proportion of men is not correlated with critical sentiment (Fig. 8B; linear regression adjusted $R^{2}=0.044$; one-sided F-test $F(1,25)=2.19$, $p=0.15$) but it is positively correlated with favorable sentiment towards collaborators (Fig. 8C; slope $s=64.37$, $95\%$ CI $(38.02,90.72)$, adjusted $R^{2}=0.483$; one-sided F-test $F(1,25)=25.31$, $p<0.0001$) and towards others in general (Fig. 8C; slope $s=48.61$, $95\%$ CI $(1.22,96.00)$, adjusted $R^{2}=0.118$; one-sided F-test $F(1,25)=4.46$, $p<0.05$). Generally, these results indicate that gender interacts with both discipline and collaboration status in explaining citation sentiment.
@@ III.3 Relations of culture across countries
Although scientists live and work within an explicitly scholarly social milieu—comprised of collaborative relationships and disciplinary identities—science and its practitioners also exist within other cultural structures outside of the academy. For example, research priorities and funding levels differ significantly across countries [125, 126]. Further, social structures of prestige and marginalization impact what science is done and who gets to do it [127, 24], often with detrimental outcomes for discovery and innovation [128, 129]. Here we asked whether the well-known cultural differences in academic writing extend to citation practices, and whether the sentiment of those citations tracks cultural norms around status, hierarchy, and power. We found that countries characterized by individualism displayed greater citation sentiment (both critical and favorable) to non-collaborators, indicating a more pervasive use of affect in scientific writing. By contrast, countries characterized by collectivism displayed less citation sentiment, utilizing a more neutral writing style. We further found that countries whose people tended to accept the unequal distribution of power (as quantified by the power-distance) displayed less citation sentiment (both critical and favorable) to non-collaborators. It is possible that citation sentiment is used in science to uphold or dismantle existing structures of status, hierarchy, and dominance whereas the lack of sentiment indicates an acceptance of power and a focus on collective goals. More broadly, our results suggest that the non-scholarly cultural systems that we as scientists are embedded within play a key role in how scholars practice science individually.
III.4 Gender differences in citation sentiment
The non-scholarly systems that scientists work within exist not only at the large scale of countries but also at the small scale of identities. A key sociocultural difference exists between those people socialized (or living) as women and those socialized (or living) as men [130]. These differences arise along multiple dimensions of experience that could impact science, for example in linguistic behavior [131] and in social behavior [132]. Consistent with these general observations, gender differences have been identified in how scholars engage in the scientific process, from writing styles [133] and research areas [134] to self-promotion practices [135], self-citation [136], positive presentation of their own results [137], and use generic statements, framing an idea as broad, timeless, and universally true [133]. Here we demonstrate that those gender differences also extend to the sentiment with which scholars cite one another. We find that men cite with greater sentiment—both critical and favorable—and women cite with greater sentiment bias, being more favorable to collaborators than to non-collaborators.
Why might men cite with more affect than women? Prior work has provided evidence that men engage socially in science, more so than women; for example, men are more likely to share their papers socially, and preferentially do so with other men [138]. This out-sized social engagement in science holds both in laboratory settings and in evaluations of real-world behavior on social media [139, 135]. Broadly, it suggests that men may treat science as a small-scale society, where prior evidence marks a key interdependence between cooperation (here perhaps supported by favorable sentiment) and status hierarchy (here perhaps upheld by critical sentiment) [109]. Why might women cite with greater sentiment bias than men, assigning more favorable (and less critical) sentiment to collaborators than to non-collaborators? Prior work suggests that women, on average, tend to use more affiliative language whereas men use more aggressive language (and are more generally talkative) [140]. Those affiliative tendencies may manifest in the use of citation sentiment to strengthen in-group relations, via more favorable and less critical language. Future work could consider how these gender differences might intersect with other dimensions of identity including race, ethnicity, sexual orientation, and socioeconomic status.
III.5 Methodological Considerations
@@
Several methodological considerations are pertinent to our study. First, our sentiment analysis method captures both evaluative judgments and expressed commonalities and differences between the citing and the cited works. Sentiment can be defined in multiple ways [141], and the findings must be interpreted in the context of the definition used. Second, because we have limited data for each individual author, we cannot predict how a single citer might behave. Instead, our results characterize trends for a group or groups of people. Third, in seeking to understand how scholars cited the work of others with different sorts of affect, it was important to minimize the effect of content. Hence, we developed and employed a null model that accounts for textual (dis)similarity between citing and cited papers, and show residual sentiment estimates. Fourth, studies of culture are challenged by the complexity of the concept, and prior work has underscored the need to think about multiple dimensions and scales of culture [70, 71]. Here we acknowledge that complexity by examining cultural factors that span from small to large scales, and concretize our investigation by using specific variables related to social practices and beliefs. Fifth, prior work has found that criticism and intellectual differences in citations are rare (from $0.8\%$ [38] to $15\%$ [142]). To ensure a reliable signal, we employed a large database of $627108$ citation sentences and a powerful and sensitive sentiment detection method using a large language model. Sixth, our results are limited to the field of neuroscience, and it is not yet clear whether they are pertinent to other fields. It seems likely that there could be links to other young disciplines, that have common factors like a relative youth which could allow for the emergence of still-malleable subdisciplines, a marked interdisciplinarity drawing on older fields, and a diversity of opinions, methods, and experimental approaches [41, 42]. It would be interesting to contrast this work to other disciplines of varying ages.
@@ V.1 Citation and Author Data
We used PubMed Central and the Web of Science API to build our datasets for analyses. The PubMed Central (PMC) Open Access Subset (OA) includes millions of research articles and pre-prints alongside relevant metadata, such as article type, venue of publication, and institutional information [144]. We gathered all English language research and review articles from PMC OA relating to neuroscience that were published between $1998$ and $2023$ and that cited at least one article. In parallel, using the Web of Science, we compiled a list of journals classified under the “neurosciences” category with an impact factor of at least $3$ in the year $2022$. We manually added three more journals creating a list of $181$ journals in which the gathered articles were published; the full list is available in the Supplement. Each article in the PMC OA subset was stored in an extensible markup language (XML) file with attached tags delineating its several components. For instance, the <article-title> tag encloses an article’s title, <abstract> encloses the abstract, and <xref> encloses a citation in the body of the text. Using these and other relevant tags, we extracted each article’s title, list of authors, affiliations (at the departmental, institutional, and country levels), and all sentences containing citations to other works in the same dataset. We obtained $69.24\%$ of last authors’ h-indices from the Web of Science. We used Gender API to infer gender from the names of the last authors, requiring the underlying model to be at least $70$% confident in its prediction [145]. The resulting dataset contained $627108$ citation sentences, $383292$ unique authors, $56034$ unique last authors; from the data, $27$ departments and $23$ countries and regions were included in the analyses. Additionally, we built a collaboration network where nodes correspond to authors and edges correspond to co-authorship. From this network, we define collaboration distance as the length of the shortest path between two individuals.
@@ V.2 Measuring Sentiment
We assessed citation sentiment using GPT-3.5-Turbo-1106, a large language model (LLM) by OpenAI. LLMs are trained on a large corpus of text data and can generate language across various domains. Through their exposure during training to diverse contexts, they can recognize subjective elements of human language, such as tone [146], irony [147, 148], sarcasm [149], and implied meaning [150]. Here, we prompted GPT-3.5-Turbo-1106 to evaluate a citation as negative, neutral, or positive; prompt text supplied to the LLM is available in the Supplement. Each entry in the citation data contained a single sentence with possibly multiple citations. Consider, for instance, the following sentence with two citations: “Lorem ipsum odor amet [citation 1], consectetuer adipiscing elit [citation 2].” This text results in two entries for the LLM to judge: (1) “Lorem ipsum odor amet [citation 1], consectetuer adipiscing elit”, and (2) “Lorem ipsum odor amet, consectetuer adipiscing elit [citation 2].” Using the OpenAI API, we passed each entry to the model separately for sentiment analysis. In practice, each citation location in a sentence was replaced by the Unicode character with encoding U+272A (see Supplement for full model prompt). It is common for an article to cite another article multiple times in different locations in the text. We aggregated overall sentiment for a given citer-citee pair through the following order of precedence: negative sentiment took precedence over positive sentiment, and both took precedence over neutral sentiment. Here, we assumed negative sentiment from a citing article overrides any positive or neutral citations [151, 152]. We used aggregated sentiment throughout the analysis.
@@
However, before using the model in this manner, we performed an additional fine-tuning step. LLMs can perform various tasks through zero-shot prompting, meaning they do not generally require more training. Supplying a well-chosen prompt is often sufficient. However, a small amount of domain-specific fine-tuning can improve performance and help generate well-formatted outputs for easier parsing. The fine-tuning process we used entailed adapting a large language model trained for a general use case to a task-specific use case, in this case, assessing citation sentiment. With the help of $5$ volunteer researchers, we manually annotated $300$ randomly chosen citations as negative ($-1$), neutral (00), or positive ($1$). These citations were drawn from journals in neuroscience and physics, the former high in critical sentiment and the latter low in critical sentiment [153], covering a broad range of sentiment expression. The fine-tuning dataset comprised these citations and their annotator-assigned labels. If the average sentiment value was less than or equal to $-0.4$, we assigned a negative label; if it is greater than or equal to $0.4$, we assigned a positive label. All citations with values between the two were assumed to be neutral. Fine-tuning GPT-3.5-Turbo-1106 with this data resulted in an improvement of $28\%$ over the baseline (see Supplement).
@@
We used a null model to benchmark observed sentiment values against those that might emerge by random chance. For instance, we found that papers with similar content were likely to involve negative citations, thereby increasing the likelihood that such a citation in the data was an outcome of the similarity of work and not an underlying bias of interest. Similar effects were associated with article type (review vs. research) and the number of times an article references another. A review article tends to be more neutral in sentiment compared to a research article, and an article that cites another article more frequently tends to be more negative. We used these factors to measure the expected probability of observing a negative, neutral, or positive sentiment for any given citer-citee pair. First, relating to shared content, we created three bins for high, medium, and low similarity, where similarity was measured as the cosine distance between the vector embeddings of article titles. The article titles were extracted using a PMC parser [154]. The embeddings were generated using OpenAI’s text-embedding-3-small model. Additionally, $15$ bins were associated with citation frequency and $2$ with article type. Initially, for each of the $90$ ($3\times 15\times 2$) resulting combinations, we performed maximum likelihood estimation to compute the probability that a citation in the data belonged to a given group. However, if a bin had fewer than $500$ samples, we used the marginal distribution over title similarity to create a larger bin. In these instances, we chose to ignore the impact of title similarity since it had the weakest effect on citation valence compared to article type and citation frequency. Further, when a citing-cited pair had a citation frequency greater than $15$, we used the corresponding bin for $15$ citations instead. We used the probabilities underlying the $54$ resulting bins to draw realizations of null sentiment values for given citing-cited pairs. These realizations served as the basis of comparison throughout our analyses.
@@
Consider a set of citations corresponding to a broader category of interest, such as citations between collaborators or citations between authors with an h-index difference of $2$. Given such a set, we can measure the ratio of positive, neutral, and negative citations within it. However, if the set is small in size, the estimated ratios may not accurately represent the true underlying distribution of sentiments for the group. To mitigate this concern, we applied bootstrapping. We created $1000$ resampled sets by randomly sampling from the original, each time with replacement, thereby allowing us to estimate a distribution of ratios for each type of sentiment. Bootstrapping in this manner helps to construct robust distributions and derive more accurate measures of variability. Simultaneously, for each of the $1000$ resampled sets for a given sentiment type, we evaluated the expected ratios considering the null model. Our measure of interest was the percentage difference in the empirical ratio relative to the ratio derived from the null model. In all figures, we plot mean values for the normalized empirical ratios and estimate error bars using standard deviations.